OpenStudy (anonymous):

You have just graduated from college and landed your first big job. You have always dreamed of being a homeowner, and after carefully shopping for your dream home, you find one that you would like to purchase at a cost of $250,000. After researching banks to find the best interest rate, you find that Banks for Homeowners offers the best rate of 6% interest that compounds monthly for 30 years. What is the monthly payment for this loan? What is the unpaid balance of the loan at the end of 5 years? What is the unpaid balance at the end of the 10th year?

OpenStudy (amistre64):

A{n} = (201/200) A{n-1} ; A{0} = 250,000

OpenStudy (amistre64):

A{0} = 250,000 (1.005)^0 A{1} = 250,000 (1.005)^1 A{2} = 250,000 (1.005)^2 A{n} = 250,000 (1.005)^n

OpenStudy (amistre64):

the interest alone is: I{n} = 250,000 (.005)^n so take how much youve spent and minus the interest to get balance left

OpenStudy (amistre64):

im think mnthly payments are determined by: 250,000 (1.005)^(360) -------------------- 30*12 but i cant be sure

OpenStudy (amistre64):

oh, and be sure to convert years to months to use in the "n spots above

OpenStudy (anonymous):

just use a fin calc for goodness sake...